# Integrating Spatiotemporal Features in LSTM for Spatially Informed COVID-19 Hospitalization Forecasting

### v1

@@ 2.4 Social-Network-Derived Auxiliary Signals
Despite these successes, social-media data have well-documented drawbacks. First, they suffer from demographic and geographic selection biases (Zhao et al. 2022). Second, they are highly vulnerable to health-related misinformation, which can distort prevalence signals and even amplify epidemic spread (Ferrara et al. 2020, DeVerna et al. 2025). Finally, operational pipelines require continuous API access and natural language processing (NLP) workflows. These dependencies might introduce processing delays and break when platform policies change—constraints incompatible with the Hub’s one-day window from data release to forecast submission. Nevertheless, Facebook’s extensive user base (with more than 2.5 billion active users globally and almost 200 million users in the U.S.), provides one of the most representative measures for broader social connectivity of spatial units, having strong potential for enhancing epidemiological forecasting.
@@ 2.5 Distinction from Existing Work
Common measures of spatial autocorrelation assign weights to each region solely by geographic contiguity or distance, which only captures local spread and ignores non-local transmission pathways (Anselin 2005, Zhang et al. 2019). Further, they assumed equal contribution from spatial neighbors, without quantifying the strength of the connection, to avoid introducing arbitrary structure into data. One of the distinctive highlights of our work is that it accounts for spatial spread by using Facebook’s Social Connectedness Index (SCI) to engineer spatiotemporal features. This index which reflects the strength of relationships between regions in the social network space, has been shown to be associated with geographic proximity, historical ties, political boundaries, and other factors (Bailey et al. 2018). Social media-derived spatial connectivity features have demonstrated a strong positive association with population movement flows, and can serve as a proxy for average human movement measurements and population interaction (Li et al. 2021). This is also supported by earlier research (Kuchler et al. 2022), showing a positive correlation between COVID-19 case clusters and SCI.
@@
While prior studies (Vahedi et al. 2021, Lucas et al. 2023) have incorporated SCI-derived spatial features for county-level case forecasting, they do not present an ablation study and they do not address the unique challenges of state-level hospitalization forecasting.
@@
Unlike previous COVID-19 case forecasting efforts, the COVID-19 Forecast Hub explicitly mandated daily, state-level, 28-day-ahead hospitalization forecasts, updated on a weekly schedule, throughout the pandemic. This is not a hypothetical exercise but an operational requirement with a core scientific challenge, as hospitalizations are considered a more reliable and policy-relevant target than cases—especially after the rise of at-home testing. Compared to county-level, weekly case forecasts, this setup has fewer total data points (51 states) yet requires forecasting at higher temporal resolution (daily rather than weekly). As a result, models must handle longer time-series with fewer spatial samples, making the learning task significantly more challenging than the discontinued case forecasting.
@@
Our work is the first to leverage social-media-derived connectivity for daily, state-level hospitalization forecasts within a deep learning framework. While a few studies incorporated SCI for county-level case predictions (Vahedi et al. 2021, Lucas et al. 2023), no prior work had harnessed these features for a 28-day operational hospitalization forecast. We address this gap by introducing our “SPH” feature to capture spatial spillover effects and integrating it into a parallel-stream LSTM architecture that learns both short- and long-term temporal dependencies. In detailed ablation experiments, we demonstrate how SPH significantly improves predictive accuracy and reduces variance—highlighting its added value in real-time policy contexts.
@@
LSTMs, a type of Recurrent Neural Networks (RNNs), excel in learning long-term temporal dependencies, making them ideal for sequence and time series data (Hochreiter and Schmidhuber 1997). LSTM and its variants have been proven successful in applications of time-series epidemiological modeling (Venna et al. 2018), especially COVID-19 Forecasting (Lucas et al. 2023, Chimmula and Zhang 2020, Shahid et al. 2020). Our model extends and improves this capability by incorporating parallel temporal lags in model input and multi-horizon ensembling strategy to further improve time series forecasts.
@@ 3 Materials and Methods
This section delineates the intricacies of data processing, feature engineering, and model architecture. We explain how we enhance our model’s predictive capacity and robustness by leveraging features derived from Facebook’s Social Connectedness Index, integrating them into a novel parallel architecture of stacked Long Short-Term Memory (LSTM) networks. Additionally, we detail our novel multi-horizon ensemble strategy aimed at preserving both the consistency and accuracy of 28-day incident hospitalization forecasting.
@@ 3.1 Data and Processing
Our forecasting target, aligned with the U.S. Forecast Hub (Cramer et al. 2022a), is the 28-day incident hospitalizations across 50 states of the U.S. plus Washington D.C. Hence, on forecast date $t$, our predictions encompass values $y_{t+1},y_{t+2},\dots,y_{t+28}$, for each of the 51 locations. We designed an autoregressive model with three input feature time series: two temporal features derived from raw incident cases and hospitalizations, respectively, and a spatiotemporal feature representing the spatial spread of the virus (and its hospitalizations) across states derived from the Facebook Social Connectedness Index.
Incident case data are sourced from the U.S. COVID-19 Forecast Hub, and the incident cases at the state-level are calculated by summing incident counts across all sub-level (county-level) locations. State-level hospitalization data are retrieved from the Department of Health and Human Services website as the sum of confirmed daily adult and pediatric COVID-19 admissions.
We use social media connectivity to capture spatial dependence. The Social Connectedness Index (SCI) is a public dataset published by Facebook and available in multiple spatial resolutions (e.g., counties, states, and countries) (Bailey et al. 2017). The index measures the strength of connectedness between two geographic areas as measured by Facebook friendship ties, providing a macro-level insight into social dynamics and disease spread potential rather than individual behaviors. Furthermore, Facebook’s extensive user base (with more than 2.5 billion active users globally and almost 200 million users in the U.S.), provides one of the most representative measures for broader social connectivity of spatial units. The SCI is calculated as (Bailey et al. 2018):
@@
where $FB\_Connections_{i,j}$ is the total number of Facebook friendship connections between users in location $i$ and users in location $j$. Similarly, $FB\_Users_{i}$ and $FB\_Users_{j}$ represent the total number of Facebook users in locations $i$ and $j$, respectively. Facebook determines geographic location based on user profiles and connection locations. Because the numerator counts mutual friendships and the denominator is the product of user counts in the two areas, SCI is mathematically symmetric, i.e. $SC_{i,j}=SC_{j,i}$. The final index is scaled to a range between 1 to 1,000,000,000. The state-level dataset covers 50 states and Washington D.C. in the U.S., which means each state has 50 values of SCIs representing the strength of connections to other states. It is worth reiterating that these connections include but are not limited to geographic proximity. The dataset is available for other countries as well, providing a valuable source for similar modeling in other regions of the world, including data-poor regions. In its current release, SCI provides connectedness indices for thousands of sub-national units across more than 200 jurisdictions worldwide, a spatial coverage unmatched by commercial cell-phone mobility products (Ilin et al. 2021, Bailey et al. 2018).
Kuchler et al. (2022) introduced a metric termed Social Proximity to Cases (SPC) to quantify the impact of social connections on disease spread. This metric captures the evolving spatial characteristics of the epidemic through a weighted sum of cases in connected counties, with weights determined by SCI. In pursuit of our goal of forecasting incident hospitalizations, we extend upon SPC to define Social Proximity to Hospitalizations (SPH). SPH represents a weighted sum of hospitalizations in connected states, encapsulating the spatial spill-over effect, i.e., the transmission of infectious disease and resulting hospitalizations from connected states. Specifically, the SPH is calculated as:
@@
where $Hospitalization\_rate_{j,t}$ represents the number of hospitalizations caused by COVID-19 per 10,000 people in state $j$ at time $t$. $SC_{i,j}$ denotes the SCI between state $i$ and $j$. $C$ is the set of all states connected to state $i$. In Eq. (4), we first row-normalize SCI for each focal state $i$. This normalization converts the symmetric SCI matrix into a directional weight vector that represents the proportion of social connections from state $i$ to every other state $j$. Consequently, SPH can be interpreted as the weighted sum of hospitalization rates in other states, weighted by how socially connected those states are to the target state $i$. While SCI remains a static annual index, SPH captures pandemic dynamics by integrating daily updated hospitalization data at time $t$. Time-series SPH values are computed for each state, alongside incident cases and hospitalizations, and are updated at each time step to effectively capture both temporal and spatial changes.
Our approach differs from most others in that we first convert raw numbers into rates (incident cases or hospitalizations divided by state population), then rescale those rates to the [-1,1] range using a MinMax scaler before feeding them into the model. Compared to raw numbers, rates are less skewed (see Fig.1), enabling the model to learn more uniformly from all states of varying populations instead of focusing on states with larger populations. Our exploratory experiments showed better predictive performance using rates over raw incidence, and more importantly, better leveraging of the SPH features regardless of state population size.
Figure 1: Histogram and Skewness Index of Raw Incident Hospitalizations and Hospitalization Rates by March 12, 2022, for all 51 states (around our evaluation period of Omicron surge). Rates are less skewed, allowing for better capturing of spatial spillover, independent of population size.
@@ 3.2 Model Architecture and Hyperparameters
We develop a parallel stacked Long Short-Term Memory (LSTM) network with spatial features to forecast multi-horizon COVID-19 hospitalizations, which we call SLSTM. Learning multi-scale temporal dependencies is essential for effectively modeling the dynamics of the pandemic, which can vary across different phases and variants. The parallel-stream design captures both short-term (7-day) trends and long-term (28-day) patterns, allowing the model to balance these temporal scales and adapt to the dynamic nature of hospitalization trends. Preliminary tests with various subperiod combinations (e.g., 7 & 14 days, 7 & 21 days, or 14 & 28 days) revealed that pairing 7 and 28 days yielded the best performance. The parallel-stream design incorporates two architecturally-identical stacked LSTM networks, with each network containing five hidden layers, as depicted in Fig.2. These networks comprise four LSTM layers with decreasing neuron counts (256, 128, 128, and 128) and a final dense layer with 64 neurons. These networks are designed to capture both short-term (7-day) trends and long-term (28-day) patterns by processing 3-channel multivariate time series arrays with different temporal lag lengths. This design allows the model to balance temporal scales and adapt to dynamic hospitalization trends by fusing short-term and long-term information. To balance short- and long-term temporal dependencies, we introduce a learnable weight, $W$, which is initialized to 1.0 and optimized via backpropagation. This weight modulates the short-term embedding before concatenation with the long-term embedding:
@@ 3.4 Multi-Horizon Ensembling Strategy
There are four main strategies for multi-step time series forecasting: direct multi-step, recursive multi-step, direct-recursive hybrid, and multiple output strategy. The direct multi-step strategy trains a separate model for each forecast time step, which is computationally burdensome and tedious to maintain. The recursive multi-step strategy only trains one model and uses its predictions as inputs to predict subsequent time steps. This is less computationally demanding than the direct strategy; however, this method propagates the errors along the time steps (Marcellino et al. 2006). In the direct-recursive hybrid strategy, a separate model is trained for each time step to be predicted, but each model may also use the predictions from models predicting prior time steps. The final strategy, the multiple output strategy, predicts the entire forecast sequence (i.e., all output time steps) at once. Multiple output models are more complex, which means they are harder to converge; however, their forecasts are more consistent across time-steps and have smaller total errors. As the multi-output models are optimized on the entire sequence, the larger time-span predictions in the sequence may have larger errors than other strategies, where individual time-step errors are minimized separately, but more consistent across the entire time-span. Lastly, all deep learning models are trained stochastically (using variations of stochastic gradient descent), resulting in varying predictions for models trained with different initial random seeds.
Our SLSTM uses a multiple-output strategy. To counter the disadvantages of the multi-output strategy, i.e., to reduce the error and variability of stochastic predictions while maintaining the consistency of different output sequences, we devise a multi-horizon ensemble strategy as follows. We design the ensemble such that for each 28 time-step in the forecast sequence, 15 different models generate a forecast for each quantile and point prediction (see Fig. 3).
@@
Among 15 prediction members for each time step, 4 models have a 7-day sequence forecast horizon, with each of those 4 having a different random initial seed. Referring to Fig. 3, it is worth noting that we stagger four subgroups of such 7-day forecast output models such that the first subgroup forecasts hospitalizations of $t+1,\dots,t+7$; the second subgroup forecasts the next 7 days of $t+8,\dots,t+14$; the third subgroup forecasts the next 7 days of $t+15,\dots,t+21$; the fourth subgroup forecast the last 7 days: $t+22,\dots,t+28$. These 7-day forecast models are expected to have less error over the 7-day output sequence (compared to optimizing a model on 28 days); however, the forecasts made by subsequent subgroups may suffer inconsistency, e.g., the difference in predictions between $t+7$ from the first subgroup and $t+8$ from the second subgroup might be large.
Therefore, we also include models with larger output sequences in the ensemble. Similarly, two staggered subgroups with seven members each are trained to forecast 14-day sequences, with seven predicting $t+1,\dots,t+14$, and seven predicting $t+15,\dots,t+28$.
Lastly, we also included four members with a direct multi-output 28-day forecast sequence. This approach allows us to generate forecasts for the entire 28-day period in a single step to keep consistency.

### v2

@@ 2.4 Social-Network-Derived Auxiliary Signals
Despite these successes, social-media data have well-documented drawbacks. First, they suffer from demographic and geographic selection biases (Zhao et al. 2022). Second, they are highly vulnerable to health-related misinformation, which can distort prevalence signals and even amplify epidemic spread (Ferrara et al. 2020, DeVerna et al. 2025). Finally, operational pipelines require continuous API access and natural language processing (NLP) workflows. These dependencies may introduce processing delays and break when platform policies change—constraints incompatible with the Hub’s one-day window from data release to forecast submission. Nevertheless, Facebook’s extensive user base (with more than 2.5 billion active users globally and almost 200 million users in the U.S.) provides one of the most representative measures of broader social connectivity of spatial units, offering strong potential for improving epidemiological forecasting.
@@ 2.5 Distinction from Existing Work
Common measures of spatial autocorrelation assign weights to each region solely by geographic contiguity or distance, which only captures local spread and ignores non-local transmission pathways (Anselin 2005, Zhang et al. 2019). Furthermore, they assume equal contributions from spatial neighbors, without quantifying the strength of the connection, to avoid introducing arbitrary structure into data. One of the distinctive highlights of our work is that it accounts for spatial spread by using Facebook’s Social Connectedness Index (SCI) to engineer spatiotemporal features. This index, which reflects the strength of relationships between regions in the social network space, has been shown to be associated with geographic proximity, historical ties, political boundaries, and other factors (Bailey et al. 2018). Social media-derived spatial connectivity features have demonstrated a strong positive association with population movement flows, and can serve as a proxy for average human movement measurements and population interaction (Li et al. 2021). This is also supported by earlier research (Kuchler et al. 2022), showing a positive correlation between COVID-19 case clusters and SCI.
@@
While prior studies (Vahedi et al. 2021, Lucas et al. 2023) have incorporated SCI-derived spatial features for county-level case forecasting, they did not present an ablation study, and they did not address the unique challenges of state-level hospitalization forecasting.
@@
Unlike previous COVID-19 case forecasting efforts, the COVID-19 Forecast Hub explicitly mandated daily, state-level, 28-day-ahead hospitalization forecasts, updated on a weekly schedule, throughout the pandemic. This is not a hypothetical exercise but an operational requirement with a core scientific challenge, as hospitalizations are considered a more reliable and policy-relevant target than cases—especially after the rise of at-home testing. Compared with county-level weekly case forecasts, this setup has fewer total data points (51 states) yet requires forecasting at a higher temporal resolution (daily rather than weekly). As a result, models must handle longer time-series with fewer spatial samples, making the learning task significantly more challenging than the discontinued case forecasting.
@@
Our work is the first to leverage social media-derived connectivity for daily, state-level hospitalization forecasts within a deep learning framework. While a few studies incorporated SCI for county-level case predictions (Vahedi et al. 2021, Lucas et al. 2023), no prior work has harnessed these features for a 28-day operational hospitalization forecast. We address this gap by introducing our “SPH” feature to capture spatial spillover effects and integrating it into a parallel-stream LSTM architecture that learns both short- and long-term temporal dependencies. In detailed ablation experiments, we demonstrate how SPH significantly improves predictive accuracy and reduces variance—highlighting its added value in real-time policy contexts.
@@
LSTMs, a type of recurrent neural network(RNN), excel in learning long-term temporal dependencies, making them ideal for sequence and time series data (Hochreiter and Schmidhuber 1997). LSTM and its variants have been proven successful in applications of time-series epidemiological modeling (Venna et al. 2018), especially COVID-19 Forecasting (Lucas et al. 2023, Chimmula and Zhang 2020, Shahid et al. 2020). Our model extends and improves this capability by incorporating parallel temporal lags in model inputs and a multi-horizon ensembling strategy to further improve time series forecasts.
@@ 3 Materials and Methods
This section delineates the intricacies of data processing, feature engineering, and model architecture. We explain how we enhanced our model’s predictive capacity and robustness by leveraging features derived from Facebook’s Social Connectedness Index, integrating them into a novel parallel architecture of stacked Long Short-Term Memory (LSTM) networks. Additionally, we detail our novel multi-horizon ensemble strategy aimed at preserving both the consistency and accuracy of 28-day incident hospitalization forecasting.
@@ 3.1 Data and Processing
Our forecasting target, aligned with the Hub (Cramer et al. 2022a), is the 28-day incident hospitalizations across 50 states of the U.S. plus Washington, D.C. Hence, on forecast date $t$, our predictions encompass values $y_{t+1},y_{t+2},\dots,y_{t+28}$ for each of the 51 locations. We designed an autoregressive model with three input feature time series: two temporal features derived from raw incident cases and hospitalizations, respectively, and a spatiotemporal feature representing the spatial spread of the virus (and its hospitalizations) across states, derived from Facebook’s Social Connectedness Index.
Incident case data were sourced from the U.S. COVID-19 Forecast Hub, and the incident cases at the state-level were calculated by summing incident counts across all sub-level (county-level) locations. State-level hospitalization data were retrieved from the Department of Health and Human Services website as the sum of confirmed daily adult and pediatric COVID-19 admissions.
We used social media connectivity to capture spatial dependence. The Social Connectedness Index (SCI) is a public dataset published by Meta and available in multiple spatial resolutions (e.g., counties, states, and countries) (Bailey et al. 2017). The index measures the strength of connectedness between two geographic areas as measured by Facebook friendship ties, providing a macro-level insight into social dynamics and disease spread potential rather than individual behaviors. Furthermore, Facebook’s extensive user base (with more than 2.5 billion active users globally and almost 200 million users in the U.S.), provides one of the most representative measures for broader social connectivity of spatial units. The SCI is calculated as (Bailey et al. 2018):
@@
where $FB\_Connections_{i,j}$ is the total number of Facebook friendship connections between users in location $i$ and users in location $j$. Similarly, $FB\_Users_{i}$ and $FB\_Users_{j}$ represent the total number of Facebook users in locations $i$ and $j$, respectively. Facebook determines a user’s geographic location based on their profile and connection locations. Because the numerator counts mutual friendships and the denominator is the product of user counts in the two areas, SCI is mathematically symmetric, i.e. $SC_{i,j}=SC_{j,i}$. The final index is scaled to a range between 1 to 1,000,000,000. The state-level dataset covers 50 states and Washington, D.C. in the U.S., which means each state has 50 values of SCIs representing the strength of connections to other states. It is worth reiterating that these connections include, but are not limited to, geographic proximity. The dataset is available for other countries as well, providing a valuable source for similar modeling in other regions of the world, including data-poor regions. In its current release, SCI provides connectedness indices for thousands of sub-national units across more than 200 jurisdictions worldwide, a spatial coverage unmatched by commercial cell-phone mobility products (Ilin et al. 2021, Bailey et al. 2018).
Kuchler et al. (2022) introduced a metric termed Social Proximity to Cases (SPC) to quantify the impact of social connections on disease spread. This metric captures the evolving spatial characteristics of the epidemic through a weighted sum of cases in connected counties, with weights determined by the SCI. In pursuit of our goal of forecasting incident hospitalizations, we extend upon SPC to define Social Proximity to Hospitalizations (SPH). SPH represents a weighted sum of hospitalizations in connected states, encapsulating the spatial spill-over effect, i.e., the transmission of infectious disease and resulting hospitalizations from connected states. Specifically, the SPH is calculated as:
@@
where $Hospitalization\_rate_{j,t}$ represents the number of hospitalizations caused by COVID-19 per 10,000 people in state $j$ at time $t$. $SC_{i,j}$ denotes the SCI between state $i$ and $j$. $C$ is the set of all states connected to state $i$. In Eq. (4), we first row-normalized the SCI for each focal state $i$. This normalization converts the symmetric SCI matrix into a directional weight vector that represents the proportion of social connections from state $i$ to every other state $j$. Consequently, SPH can be interpreted as the weighted sum of hospitalization rates in other states, weighted by how socially connected those states are to the target state $i$. While SCI remains a static annual index, SPH captures pandemic dynamics by integrating daily updated hospitalization data at time $t$. Time-series SPH values are computed for each state, alongside incident cases and hospitalizations, and are updated at each time step to effectively capture both temporal and spatial changes.
Our approach differed from most others in that we first converted raw numbers into rates (incident cases or hospitalizations divided by state population), then rescaled those rates to the [-1,1] range using a MinMax scaler before feeding them into the model. Compared to raw numbers, rates are less skewed (see Fig.1), enabling the model to learn more uniformly from all states of varying populations instead of focusing on states with larger populations. Our exploratory experiments showed better predictive performance using rates over raw incidence, and more importantly, better leveraging of the SPH features regardless of state population size.
Figure 1: Histogram and Skewness Index of Raw Incident Hospitalizations and Hospitalization Rates by March 12, 2022, for all 51 states (around our evaluation period of the Omicron surge). Rates are less skewed, allowing for a better capture of spatial spillovers, independent of population size.
@@ 3.2 Model Architecture and Hyperparameters
We developed a parallel stacked Long Short-Term Memory (LSTM) network with spatial features to forecast multi-horizon COVID-19 hospitalizations, which we refer to as SLSTM. Learning multi-scale temporal dependencies is crucial for effectively modeling the dynamics of the pandemic, which can vary across different phases and variants. The parallel-stream design captures both short-term (7-day) trends and long-term (28-day) patterns, allowing the model to balance these temporal scales and adapt to the dynamic nature of hospitalization trends. Preliminary tests with various subperiod combinations (e.g., 7 & 14 days, 7 & 21 days, or 14 & 28 days) revealed that pairing 7 and 28 days yielded the best performance. The parallel-stream design incorporates two architecturally-identical stacked LSTM networks, with each network containing five hidden layers, as depicted in Fig.2. These networks comprise four LSTM layers with decreasing neuron counts (256, 128, 128, and 128) and a final dense layer with 64 neurons. These networks are designed to capture both short-term (7-day) trends and long-term (28-day) patterns by processing 3-channel multivariate time series arrays with different temporal lag lengths. This design allows the model to balance temporal scales and adapt to dynamic hospitalization trends by fusing short-term and long-term information. To balance short- and long-term temporal dependencies, we introduced a learnable weight, $W$, which was initialized to 1.0 and optimized via backpropagation. This weight modulates the short-term embedding before concatenation with the long-term embedding:
@@ 3.4 Multi-Horizon Ensembling Strategy
There are four main strategies for multi-step time series forecasting: direct multi-step, recursive multi-step, direct-recursive hybrid, and multiple output strategy. The direct multi-step strategy trains a separate model for each forecast time step, which is computationally burdensome and tedious to maintain. The recursive multi-step strategy only trains one model and uses its predictions as inputs to predict subsequent time steps. This is less computationally demanding than the direct strategy; however, this approach propagates errors along the time steps (Marcellino et al. 2006). In the direct-recursive hybrid strategy, a separate model is trained for each time step to be predicted, but each model may also use the predictions from models predicting prior time steps. The final strategy, the multiple output strategy, predicts the entire forecast sequence (i.e., all output time steps) at once. Multiple-output models are more complex, which means they are harder to converge; however, their forecasts are more consistent across time steps and have smaller total errors. As the multi-output models are optimized over the entire sequence, the larger time-span predictions in the sequence may have larger errors than other strategies, where individual time-step errors are minimized separately, but more consistent across the entire time-span. Lastly, all deep learning models are trained stochastically (using variations of stochastic gradient descent), resulting in varying predictions for models trained with different initial random seeds.
Our SLSTM uses a multiple-output strategy. To counter the disadvantages of the multi-output strategy, i.e., to reduce the error and variability of stochastic predictions while maintaining the consistency of different output sequences, we devised a multi-horizon ensemble strategy as follows. We designed the ensemble such that for each 28 time-step in the forecast sequence, 15 different models generate a forecast for each quantile and point prediction (see Fig. 3).
@@
Among the 15 prediction members for each time step, 4 models have a 7-day sequence forecast horizon, with each of those 4 having a different random initial seed. Referring to Fig. 3, it is worth noting that we staggered four subgroups of such 7-day forecast output models such that the first subgroup forecasts hospitalizations of $t+1,\dots,t+7$; the second subgroup forecasts the next 7 days of $t+8,\dots,t+14$; the third subgroup forecasts the next 7 days of $t+15,\dots,t+21$; the fourth subgroup forecast the last 7 days: $t+22,\dots,t+28$. These 7-day forecast models are expected to have less error over the 7-day output sequence (compared to optimizing a model on 28 days); however, the forecasts made by subsequent subgroups may suffer inconsistency, e.g., the difference in predictions between $t+7$ from the first subgroup and $t+8$ from the second subgroup might be large.
Therefore, we also included models with larger output sequences in the ensemble. Similarly, two staggered subgroups with seven members each are trained to forecast 14-day sequences, with seven members predicting $t+1,\dots,t+14$, and seven members predicting $t+15,\dots,t+28$.
Lastly, we also included four members with a direct multi-output 28-day forecast sequence. This approach enables us to generate forecasts for the entire 28-day period in a single step, thereby maintaining consistency.